# It’s always the quiet ones: Single Degenerate Double Detonation Type Ia Supernova from Quiescent Helium Accretion

### v1

@@ Abstract
1 Introduction
Type Ia supernovae (SNe Ia) are indispensable tools for cosmology (Brout et al., 2019; Galbany et al., 2023; Palicio et al., 2024) and major contributors to galactic iron-peak enrichment (Tang & Wang, 2010; De et al., 2014; Cavichia et al., 2024; Bhattacharya et al., 2025), yet decades of study have not yielded a consensus on their progenitors (Howell, 2011; Maoz et al., 2014; Hoeflich, 2017; Jha et al., 2019; Liu et al., 2023; Soker, 2024, 2025; Ruiter & Seitenzahl, 2025). Classical single-degenerate (SD) models (Whelan & Iben, 1973; Nomoto, 1982a) invoke a carbon–oxygen white dwarf (CO WD) accreting from a non-degenerate companion until reaching the Chandrasekhar mass ($\sim\!1.4\,\mathrm{M_{\odot}}$), where ignition proceeds via a deflagration-to-detonation transition (Khokhlov, 1991; Plewa et al., 2004; Gamezo et al., 2004, 2005) or a gravitationally confined detonation in multi-dimensional simulations (Calder et al., 2004; Meakin et al., 2009; Jordan et al., 2008, 2012). However, deep late-time nebular spectra of nearby SNe Ia (e.g., SN 2011fe) show no $\mathrm{H}\alpha$ or He i emission down to $\lesssim\!10^{-4}\,\mathrm{M_{\odot}}$ of stripped companion material, and pre-explosion imaging together with early-time UV limits rule out a bright red-giant, placing stringent constraints on classic SD channels (Li et al., 2011; Nugent et al., 2011). Observations of SN 2022xkq and SN 2023bee—showing fast declines and/or early flux excesses—pose additional challenges to current progenitor and explosion models (Pearson et al., 2024; Wang et al., 2024).
An alternative double-degenerate (DD) pathway involves the merger of two WDs via gravitational-wave–driven inspiral. If the combined mass exceeds $\sim\!1.4\,\mathrm{M_{\odot}}$, central carbon ignition may occur (“classical” DD; Webbink 1984; Iben & Tutukov 1984), but off-center burning can instead form an O–Ne WD rather than a supernova (Saio & Nomoto, 1998; Shen et al., 2012). More recently, “dynamically driven double-degenerate double detonations” ($\mathrm{D}^{6}$; Shen et al. 2010; Pakmor et al. 2013; Boos et al. 2024; Pollin et al. 2025) and head-on collisions (Raskin et al., 2009; Kushnir et al., 2013; Glanz et al., 2025) have been shown to produce prompt, sub-Chandrasekhar detonations in one or both WDs. Both SD and DD channels predict delay-time distributions (DTD) $\propto t^{-1}$ over 0.10.11010 Gyr (Friedmann & Maoz, 2018; Freundlich & Maoz, 2021), consistent with observed SN Ia rates in old stellar populations, yet neither channel cleanly satisfies all spectroscopic and rate constraints.
@@ 1 Introduction
Against this backdrop, the sub-Chandrasekhar “double-detonation” scenario has emerged as a robust alternative: a CO WD accretes a thin He shell that detonates first, sending a shock into the core that triggers a secondary carbon–oxygen detonation (Nomoto, 1982b). Multi-dimensional studies demonstrate that even shells as low as $\sim\!0.03\,\mathrm{M_{\odot}}$ can focus a strong inward shock and robustly detonate the core—yielding normal-luminosity SNe Ia without fine tuning (Pakmor et al., 2022). Synthetic light curves predict early UV/optical “bumps” from shell ashes, and high-velocity Ca ii/Ti ii absorption features in the first days—features that are now being sought in high-cadence surveys (Yao et al., 2019; Piro & Nakar, 2025).
A critical uncertainty in these models has been the physical origin and composition of the He shell. One scenario that in certain cases may allow for a helium shell to gradually accumulate on the surface of a WD, is as the product of nova eruptions. In this case, the system would be a cataclysmic variable (CV) where mass in transferred through the inner Lagrangian point (L1) Shara (1981); Prialnik & Kovetz (1995); Yaron et al. (2005); Shara et al. (2018); Hachisu & Kato (2010), or a symbiotic system, for which the wind from a red giant (RG) or an asymptotic giant branch star (AGB) is either gravitationally captured by the WD via the Bondi-Hoyle-Littleton (BHL) mechanism, Mikolajewska & Kenyon (1992); Mikolajewska (2008, 2010); Hillman & Kashi (2021); Vathachira et al. (2024, 2025a) or if the binary orbit is close enough, via wind focusing through L1, i.e., wind Roche-lobe overflow (WRLOF) Abate et al. (2013); Saladino et al. (2019); Vathachira et al. (2025b). Regardless of the mass transfer mechanism, the result is hydrogen-rich material being accumulated on the surface of the degenerate WD, which eventually ignites in a runaway process and produces a nova eruption (e.g., Gallagher & Starrfield, 1978; Starrfield & Sparks, 1987; Warner, 1995). Models have shown that if the hydrogen (H) is accreted at a high enough rate, the eruption will not eject all the mass that has been accreted, but will leave behind a He-rich residue Kato & Hachisu (1999, 2004); Hillman et al. (2016). This helium may accumulate over many H-novae, until eventually reaching its own critical mass that triggers a He-nova José et al. (1993); Idan et al. (2013); Newsham et al. (2014). Based on this, we Hillman et al. (2015) followed the results of the one-dimensional nova-simulation code accreting hydrogen at a high rate, and used the average accumulation rate of He-rich material as the input for direct helium accretion. Our results showed periodic helium novae that, after a long adaptation period, became less degenerate and led to the growth of the CO core.
@@
1.1 Double-detonations with thick vs. thin helium shells
Early sub-Chandrasekhar studies showed that detonating a sufficiently massive helium (He) shell can drive a converging shock into the C/O core and ignite a secondary detonation (e.g., Livne, 1990; Woosley & Weaver, 1994). These “thick-shell” models (typically $M_{\rm He}\!\gtrsim\!0.1~M_{\odot}$) also highlighted a tension: extensive He-shell burning at high velocity seeds outer iron-group and intermediate-mass material (notably Ti, Cr, Ca) that can produce strong line blanketing and redder colors than normal SNe Ia (Woosley & Weaver, 1994). Motivated by this, the field pivoted to “thin-shell” realizations where $M_{\rm He}\!\lesssim\!0.05~M_{\odot}$ suffices to trigger the core while minimizing high-velocity shell ash (Fink et al., 2010; Kromer et al., 2010; Woosley & Kasen, 2011).
@@
1.2 Relation to prior simulations
@@
Multi-D work since the 1990s finds that once a lateral He detonation wraps the star, the inward shock can reach core-ignition conditions over broad parameters (Livne, 1990; Fink et al., 2010). Where models diverge is in the observable imprint of the shell ash. Thin, modestly polluted shells ($M_{\rm He}\!\lesssim\!0.05~M_{\odot}$) tend to yield bluer colors and spectra closer to normal SNe Ia (Kromer et al., 2010; Woosley & Kasen, 2011), whereas thicker shells enhance Ca/Ti/Cr at high velocity and can redden the early-time SED (Woosley & Kasen, 2011). Systematic grids connecting shell mass/composition to light-curve and spectral behavior (Polin et al., 2019) further indicate that the “massive-shell” regime better matches peculiar, strongly line-blanketed cases (e.g., SN 2018byg) than the bulk of normal SNe Ia.
@@
Our fiducial model lies on the “thicker-shell” side of these mappings, but with a stratification and small residual He mass that may mitigate the most extreme color effects (see Section 4.1.1). Here we emphasize the dynamical robustness and nucleosynthetic budgets and quantify the outer-shell IME/IGE yields against the literature ranges; detailed radiative transfer is beyond this paper’s scope.
2 Methods
@@
We begin from the quiescent helium-accretion model #23 of Hillman et al. (2025) (see there, Tables 1 and 3). A $0.7\,\mathrm{M_{\odot}}$ carbon–oxygen (CO) white dwarf (WD) accretes helium at $\dot{M}=10^{-8}\,\mathrm{M_{\odot}\,yr^{-1}}$ and grows to $1.1\,\mathrm{M_{\odot}}$. The one–dimensional (1D) hydrodynamic Lagrangian nova code used by Hillman et al. (2025) follows consecutive eruptions from helium accretion. Originally developed for hydrogen–rich accretion (Prialnik & Kovetz, 1995; Yaron et al., 2005; Epelstain et al., 2007; Hillman et al., 2015), the code was adapted for helium accretion as in Hillman et al. (2016) to quantify how much helium a WD can retain prior to a thermonuclear runaway (TNR). We extract a snapshot immediately before explosive burning (i.e., before Si–group burning becomes numerically unresolved in 1D), yielding hydrostatic radial profiles $\{\rho(r),T(r),X_{i}(r)\}$ of a helium–rich envelope over a CO core as functions of enclosed mass and radius. Figure 1 summarizes the long–term envelope growth and core composition evolution for several $(M_{\rm WD,i},\dot{M})$ tracks from Hillman et al. (2025); our fiducial track is $M_{\rm WD,i}=0.7\,\mathrm{M_{\odot}}$, $\dot{M}=10^{-8}\,\mathrm{M_{\odot}\,yr^{-1}}$.
@@
The explosive phase is modeled with flash 4.7 (Fryxell et al., 2000), solving the compressible Euler equations in three dimensions with the unsplit solver (Lee, 2006; Lee & Deane, 2009; Lee, 2013) and self–gravity via the multipole Poisson solver. Microphysics includes the Helmholtz electron–positron Fermi–Dirac equation of state (degenerate/relativistic electrons and positrons, ideal ions with Coulomb corrections, and radiation; Aparicio (1998); Timmes & Arnett (1999); Timmes & Swesty (2000)), and nuclear energy generation with the approximate 19–isotope $\alpha$–chain network (Weaver et al., 1978), which captures the dominant energy–releasing flows for helium burning and CO detonation.
@@ 4 Discussion
Independent, remnant–phase support for this layered structure has emerged from integral–field spectroscopy of the young LMC remnant SNR 0509–67.5. High–resolution MUSE data reveal a double–shell morphology in highly ionized calcium ([Ca XV]) and a single sulphur shell in the reverse–shocked ejecta. The outer Ca shell is naturally interpreted as the product of the helium–layer detonation, while the inner Ca shell arises from incomplete burning in the CO–core detonation—precisely the stratification expected from sub–Chandrasekhar double detonations. This “double–shell Ca” fingerprint provides spatially resolved, remnant–phase evidence that the double–detonation mechanism operates in nature (Das et al., 2025).
Second, the total 56Ni mass is adequate for normal peak luminosities without requiring near-Chandrasekhar densities or strong electron captures. The comparatively low stable iron–peak content ($M(\mathrm{Fe})\simeq 0.018\,\mathrm{M_{\odot}}$; trace Mn/Cr) is consistent with sub–Chandrasekhar densities at burning and should imprint itself in nebular ratios (e.g., [Fe iii]/[Fe ii], [Ni ii]/[Fe ii]) at $\gtrsim 200$ days. Because the small off–axis temperature perturbation is rapidly overtaken by the wrapping He detonation, global asymmetries are muted at homology and are unlikely to leave observable signatures or strong continuum polarization.
@@ 4.1 Caveats
4.1.1 IME/IGE in the outer ejecta
A known vulnerability of thick-shell double detonations is the chemical makeup of the outer $v\!\gtrsim\!1.5\times 10^{4}~\mathrm{km\,s^{-1}}$ layers. He detonations at $\rho\!\sim\!10^{5}$$10^{6}~\mathrm{g\,cm^{-3}}$ can synthesize significant Ca and, depending on pollution, Ti/Cr, while leaving some residual He (Woosley & Weaver, 1994; Shen & Moore, 2014). These species shape early UV/blue opacity and can drive prominent high-velocity Ca ii features. Grids and case studies indicate that more massive shell or shells with low C/N pollution tend to produce stronger UV suppression and redder grg-r at early times, with SN 2018byg providing an extreme instance (Polin et al., 2019; De et al., 2019).
To make these risks explicit, we tabulate (i) $M(\mathrm{Ca}),M(\mathrm{Ti}),M(\mathrm{Cr})$ above $1.5\times 10^{4}~\mathrm{km\,s^{-1}}$, (ii) the residual He mass in the same velocity layer, and (iii) the angle-averaged velocity distributions of these species (Table 1 and Fig. 4). While radiative transfer is deferred, these numbers can be read against published thresholds for strong line blanketing and HVFs. We also note that modest shell pollution by C/O or 14N reduces Ti/Cr yields and can lower He i detectability by altering ionization/excitation conditions (Boyle et al., 2017; Magee et al., 2022).
5 Conclusions and Summary

### v2

@@ Abstract
I Introduction
Type Ia supernovae (SNe Ia) are indispensable tools for cosmology (Brout et al., 2019; Galbany et al., 2023; Palicio et al., 2024) and major contributors to galactic iron-peak enrichment (Tang and Wang, 2010; De et al., 2014; Cavichia et al., 2024; Bhattacharya et al., 2025), yet decades of study have not yielded a consensus on their progenitors (Howell, 2011; Maoz et al., 2014; Hoeflich, 2017; Jha et al., 2019; Liu et al., 2023; Soker, 2024, 2025; Ruiter and Seitenzahl, 2025). Classical single-degenerate (SD) models (Whelan and Iben, 1973; Nomoto, 1982a) invoke a carbon–oxygen white dwarf (CO WD) accreting from a non-degenerate companion until reaching the Chandrasekhar mass ($\sim\!1.4\,\mathrm{M_{\odot}}$), where ignition proceeds via a deflagration-to-detonation transition (Khokhlov, 1991; Plewa et al., 2004; Gamezo et al., 2004, 2005) or a gravitationally confined detonation in multi-dimensional simulations (Calder et al., 2004; Meakin et al., 2009; Jordan et al., 2008, 2012). However, deep late-time nebular spectra of nearby SNe Ia (e.g., SN 2011fe) show no $\mathrm{H}\alpha$ or He i emission down to $\lesssim\!10^{-4}\,\mathrm{M_{\odot}}$ of stripped companion material, and pre-explosion imaging together with early-time UV limits rule out a bright red-giant, placing stringent constraints on classic SD channels (Li et al., 2011; Nugent et al., 2011). Observations of SN 2022xkq and SN 2023bee—showing fast declines and/or early flux excesses—pose additional challenges to current progenitor and explosion models (Pearson et al., 2024; Wang et al., 2024).
An alternative double-degenerate (DD) pathway involves the merger of two WDs via gravitational-wave–driven inspiral. If the combined mass exceeds $\sim\!1.4\,\mathrm{M_{\odot}}$, central carbon ignition may occur (“classical” DD; Webbink 1984; Iben and Tutukov 1984), but off-center burning can instead form an O–Ne WD rather than a supernova (Saio and Nomoto, 1998; Shen et al., 2012). More recently, “dynamically driven double-degenerate double detonations” ($\mathrm{D}^{6}$; Shen et al. 2010; Pakmor et al. 2013; Boos et al. 2024; Pollin et al. 2025) and head-on collisions (Raskin et al., 2009; Kushnir et al., 2013; Glanz et al., 2025) have been shown to produce prompt, sub-Chandrasekhar detonations in one or both WDs. Both SD and DD channels predict delay-time distributions (DTD) $\propto t^{-1}$ over 0.10.11010 Gyr (Friedmann and Maoz, 2018; Freundlich and Maoz, 2021), consistent with observed SN Ia rates in old stellar populations, yet neither channel cleanly satisfies all spectroscopic and rate constraints.
@@ I Introduction
Against this backdrop, the sub-Chandrasekhar “double-detonation” scenario has emerged as a robust alternative: a CO WD accretes a thin He shell that detonates first, sending a shock into the core that triggers a secondary carbon–oxygen detonation (Nomoto, 1982b). Multi-dimensional studies demonstrate that even shells as low as $\sim\!0.03\,\mathrm{M_{\odot}}$ can focus a strong inward shock and robustly detonate the core—yielding normal-luminosity SNe Ia without fine tuning (Pakmor et al., 2022). Synthetic light curves predict early UV/optical “bumps” from shell ashes, and high-velocity Ca ii/Ti ii absorption features in the first days—features that are now being sought in high-cadence surveys (Yao et al., 2019; Piro and Nakar, 2025).
A critical uncertainty in these models has been the physical origin and composition of the He shell. One scenario that in certain cases may allow for a helium shell to gradually accumulate on the surface of a WD, is as the product of nova eruptions. In this case, the system would be a cataclysmic variable (CV) where mass in transferred through the inner Lagrangian point (L1) Shara (1981); Prialnik and Kovetz (1995); Yaron et al. (2005); Shara et al. (2018); Hachisu and Kato (2010), or a symbiotic system, for which the wind from a red giant (RG) or an asymptotic giant branch star (AGB) is either gravitationally captured by the WD via the Bondi-Hoyle-Littleton (BHL) mechanism, Mikolajewska and Kenyon (1992); Mikolajewska (2008, 2010); Hillman and Kashi (2021); Vathachira et al. (2024, 2025b) or if the binary orbit is close enough, via wind focusing through L1, i.e., wind Roche-lobe overflow (WRLOF) Abate et al. (2013); Saladino et al. (2019); Vathachira et al. (2025a). Regardless of the mass transfer mechanism, the result is hydrogen-rich material being accumulated on the surface of the degenerate WD, which eventually ignites in a runaway process and produces a nova eruption (e.g., Gallagher and Starrfield, 1978; Starrfield and Sparks, 1987; Warner, 1995). Models have shown that if the hydrogen (H) is accreted at a high enough rate, the eruption will not eject all the mass that has been accreted, but will leave behind a He-rich residue Kato and Hachisu (1999, 2004); Hillman et al. (2016). This helium may accumulate over many H-novae, until eventually reaching its own critical mass that triggers a He-nova José et al. (1993); Idan et al. (2013); Newsham et al. (2014). Based on this, we Hillman et al. (2015) followed the results of the one-dimensional nova-simulation code accreting hydrogen at a high rate, and used the average accumulation rate of He-rich material as the input for direct helium accretion. Our results showed periodic helium novae that, after a long adaptation period, became less degenerate and led to the growth of the CO core.
@@
I.1 Double-detonations with thick vs. thin helium shells
Early sub-Chandrasekhar studies showed that detonating a sufficiently massive helium (He) shell can drive a converging shock into the C/O core and ignite a secondary detonation (e.g., Livne, 1990; Woosley and Weaver, 1994). These “thick-shell” models (typically $M_{\rm He}\!\gtrsim\!0.1~M_{\odot}$) also highlighted a tension: extensive He-shell burning at high velocity seeds outer iron-group and intermediate-mass material (notably Ti, Cr, Ca) that can produce strong line blanketing and redder colors than normal SNe Ia (Woosley and Weaver, 1994). Motivated by this, the field pivoted to “thin-shell” realizations where $M_{\rm He}\!\lesssim\!0.05~M_{\odot}$ suffices to trigger the core while minimizing high-velocity shell ash (Fink et al., 2010; Kromer et al., 2010; Woosley and Kasen, 2011).
@@
I.2 Relation to prior simulations
@@
Multi-D work since the 1990s finds that once a lateral He detonation wraps the star, the inward shock can reach core-ignition conditions over broad parameters (Livne, 1990; Fink et al., 2010). Where models diverge is in the observable imprint of the shell ash. Thin, modestly polluted shells ($M_{\rm He}\!\lesssim\!0.05~M_{\odot}$) tend to yield bluer colors and spectra closer to normal SNe Ia (Kromer et al., 2010; Woosley and Kasen, 2011), whereas thicker shells enhance Ca/Ti/Cr at high velocity and can redden the early-time SED (Woosley and Kasen, 2011). Systematic grids connecting shell mass/composition to light-curve and spectral behavior (Polin et al., 2019) further indicate that the “massive-shell” regime better matches peculiar, strongly line-blanketed cases (e.g., SN 2018byg) than the bulk of normal SNe Ia.
@@
Our fiducial model lies on the “thicker-shell” side of these mappings, but with a stratification and small residual He mass that may mitigate the most extreme color effects (see Section IV.1.1). Here we emphasize the dynamical robustness and nucleosynthetic budgets and quantify the outer-shell IME/IGE yields against the literature ranges; detailed radiative transfer is beyond this paper’s scope.
II Methods
@@
We begin from the quiescent helium-accretion model #23 of Hillman et al. (2025) (see there, Tables 1 and 3). A $0.7\,\mathrm{M_{\odot}}$ carbon–oxygen (CO) white dwarf (WD) accretes helium at $\dot{M}=10^{-8}\,\mathrm{M_{\odot}\,yr^{-1}}$ and grows to $1.1\,\mathrm{M_{\odot}}$. The one–dimensional (1D) hydrodynamic Lagrangian nova code used by Hillman et al. (2025) follows consecutive eruptions from helium accretion. Originally developed for hydrogen–rich accretion (Prialnik and Kovetz, 1995; Yaron et al., 2005; Epelstain et al., 2007; Hillman et al., 2015), the code was adapted for helium accretion as in Hillman et al. (2016) to quantify how much helium a WD can retain prior to a thermonuclear runaway (TNR). We extract a snapshot immediately before explosive burning (i.e., before Si–group burning becomes numerically unresolved in 1D), yielding hydrostatic radial profiles $\{\rho(r),T(r),X_{i}(r)\}$ of a helium–rich envelope over a CO core as functions of enclosed mass and radius. Figure 1 summarizes the long–term envelope growth and core composition evolution for several $(M_{\rm WD,i},\dot{M})$ tracks from Hillman et al. (2025); our fiducial track is $M_{\rm WD,i}=0.7\,\mathrm{M_{\odot}}$, $\dot{M}=10^{-8}\,\mathrm{M_{\odot}\,yr^{-1}}$.
@@
The explosive phase is modeled with flash 4.7 (Fryxell et al., 2000), solving the compressible Euler equations in three dimensions with the unsplit solver (Lee, 2006; Lee and Deane, 2009; Lee, 2013) and self–gravity via the multipole Poisson solver. Microphysics includes the Helmholtz electron–positron Fermi–Dirac equation of state (degenerate/relativistic electrons and positrons, ideal ions with Coulomb corrections, and radiation; Aparicio (1998); Timmes and Arnett (1999); Timmes and Swesty (2000)), and nuclear energy generation with the approximate 19–isotope $\alpha$–chain network (Weaver et al., 1978), which captures the dominant energy–releasing flows for helium burning and CO detonation.
@@ IV Discussion
Independent, remnant–phase support for this layered structure has emerged from integral–field spectroscopy of the young LMC remnant SNR 0509–67.5. High–resolution MUSE data reveal a double–shell morphology in highly ionized calcium ([Ca XV]) and a single sulphur shell in the reverse–shocked ejecta. The outer Ca shell is naturally interpreted as the product of the helium–layer detonation, while the inner Ca shell arises from incomplete burning in the CO–core detonation—precisely the stratification expected from sub–Chandrasekhar double detonations. This “double–shell Ca” fingerprint provides spatially resolved, remnant–phase evidence that the double–detonation mechanism operates in nature (Das et al., 2025; Soker2025double–shell-Ca). Second, the total 56Ni mass is adequate for normal peak luminosities without requiring near-Chandrasekhar densities or strong electron captures. The comparatively low stable iron–peak content ($M(\mathrm{Fe})\simeq 0.018\,\mathrm{M_{\odot}}$; trace Mn/Cr) is consistent with sub–Chandrasekhar densities at burning and should imprint itself in nebular ratios (e.g., [Fe iii]/[Fe ii], [Ni ii]/[Fe ii]) at $\gtrsim 200$ days. Because the small off–axis temperature perturbation is rapidly overtaken by the wrapping He detonation, global asymmetries are muted at homology and are unlikely to leave observable signatures or strong continuum polarization.
@@ IV.1 Caveats
IV.1.1 IME/IGE in the outer ejecta
A known vulnerability of thick-shell double detonations is the chemical makeup of the outer $v\!\gtrsim\!1.5\times 10^{4}~\mathrm{km\,s^{-1}}$ layers. He detonations at $\rho\!\sim\!10^{5}$$10^{6}~\mathrm{g\,cm^{-3}}$ can synthesize significant Ca and, depending on pollution, Ti/Cr, while leaving some residual He (Woosley and Weaver, 1994; Shen and Moore, 2014). These species shape early UV/blue opacity and can drive prominent high-velocity Ca ii features. Grids and case studies indicate that more massive shell or shells with low C/N pollution tend to produce stronger UV suppression and redder grg-r at early times, with SN 2018byg providing an extreme instance (Polin et al., 2019; De et al., 2019).
To make these risks explicit, we tabulate (i) $M(\mathrm{Ca}),M(\mathrm{Ti}),M(\mathrm{Cr})$ above $1.5\times 10^{4}~\mathrm{km\,s^{-1}}$, (ii) the residual He mass in the same velocity layer, and (iii) the angle-averaged velocity distributions of these species (Table 1 and Fig. 4). While radiative transfer is deferred, these numbers can be read against published thresholds for strong line blanketing and HVFs. We also note that modest shell pollution by C/O or 14N reduces Ti/Cr yields and can lower He i detectability by altering ionization/excitation conditions (Boyle et al., 2017; Jiang et al., 2017; Magee et al., 2022).
V Conclusions and Summary